OpenStudy (anonymous):

this one confuses me quick helpp please 7u 14(u+v)

OpenStudy (anonymous):

7u over 14(u+v)

OpenStudy (anonymous):

confuses me too since there is no question here, just an unreadable expression

OpenStudy (anonymous):

\[\frac{7u}{14(u+v)}\] ?

OpenStudy (anonymous):

would it be u/2u+2v

OpenStudy (anonymous):

yea how do u type it in like that

myininaya (myininaya):

based on what you did, it looks like you are asked to simplify the expression?

OpenStudy (anonymous):

yes if the question is to reduce, it will be \[\frac{u}{2(u+v)}\] or \[\frac{u}{2u+2v}\]

OpenStudy (anonymous):

well yes u are correct on that

OpenStudy (anonymous):

so i was good

OpenStudy (anonymous):

thank you satellite

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yw